OpenStudy (anonymous):

x+1 over x-1 = 14 over 20

OpenStudy (anonymous):

\[(x + 1) \div (x - 1) = 14 \div 20\]

OpenStudy (anonymous):

set up as a proportion x + 1 = 14 ----- --- x - 1 20 cross multiply to get (x + 1)(20) = (x - 1)(14)

OpenStudy (anonymous):

then distribute: 20x + 20 = 14x - 14 solve for x